Circular No. 7222

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 N2 Daniel W. Lynn, Kinglake West, Victoria, Australia, reports his vi-

sual discovery of a comet using handheld 10×50 binoculars. The following observations have been reported:

1999 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 13.451  9^{h}41.^{m}0  − 16^{o}30′  8.0 Lynn
13.951  9 47.5  −15 25  7.0 Neto
14.37580 9 52 03.^{s}27 −14 29 18′′.8 8.3 Zoltowski
14.37842 9 52 05.19 −14 28 57.6  _{”}
14.38241 9 52 08.06 −14 28 26.1  _{”}
14.38673 9 52 11.21 −14 27 51.4  _{”}
14.39091 9 52 14.25 −14 27 17.8  _{”}
14.394 9 52.4 −14 25 8.0 Lynn
14.39624 9 52 18.12 −14 26 35.1  Zoltowski
14.40275 9 52 22.83 −14 25 43.1  _{”}
14.40995 9 52 28.05 −14 24 45.4  _{”}

D. W. Lynn (near Melbourne). 10×50 binoculars and 0.20-m f/6 reflector (40×, 150×). Comet low in sky, moving northeastward at ∼ 10′/hr. Coma diameter 4′.2, fairly strong central brightening, but no discernible nuclear condensation and no tail.

V. F. de Assis Neto (São Francisco de Oliveira, Minas Gerais, Brazil).

10×70 binoculars. Communicated by C. Jacques. F. B. Zoltowski (Woomera). 0.30-m f/10 Schmidt-Cassegrain reflector +

CCD. Fairly dense coma of diameter ∼ 2′; narrow tail > 20′ long in p.a. 140^{o}; m_{2} = 11.6–12.1.

COMET C/1999 N3 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports observations of another apparent Kreutz sungrazing comet that he discovered in images taken with the C3 coronagraph aboard SOHO. Astrometric measurements made by Biesecker (and reduced by B. G. Marsden), together with orbital elements by Marsden, are given on MPEC 1999-N24.

1999 UT  α_{2000}  δ_{2000}
July 9.363  7^{h}00.^{m}2  +19^{o}40′
1999 July 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green